# Revisiting In-Context Learning with Long Context Language Models

### v2

@@ Abstract
In-Context Learning (ICL) is a technique by which language models make predictions based on examples provided in their input context. Previously, their context window size imposed a limit on the number of examples that can be shown, making example selection techniques crucial for identifying the maximally effective set of examples. However, the recent advent of Long Context Language Models (LCLMs) has significantly increased the number of examples that can be included in context, raising an important question of whether ICL performance in a many-shot regime is still sensitive to the method of sample selection. To answer this, we revisit these approaches in the context of LCLMs through extensive experiments on 18 datasets spanning 4 tasks. Surprisingly, we observe that sophisticated example selection techniques do not yield significant improvements over a simple random sample selection method. Instead, we find that the advent of LCLMs has fundamentally shifted the challenge of ICL from that of selecting the most effective examples to that of collecting sufficient examples to fill the context window. Specifically, in certain datasets, including all available examples does not fully utilize the context window; however, by augmenting the examples in context with a simple data augmentation approach, we substantially improve ICL performance by 5%.
@@
Jinheon Baek1thanks: This work was conducted during an internship at Google. For correspondence, contact jinheon.baek@kaist.ac.kr.   Sun Jae Lee2   Prakhar Gupta2 Geunseob (GS) Oh2Siddharth Dalmia2Prateek Kolhar2 KAIST1    Google DeepMind2
@@ 1 Introduction
Figure 1: Results of various sample selection approaches in many-shot ICL with LCLMs. Approaches include Retrieval that selects examples similar to the target query, Diversity that aims for maximizing example variety, Curriculum that arranges examples in order from easiest to hardest, and Hard that uses only challenging examples, alongside Random that selects examples without any constraints. Results indicate that sample selection methods provide no significant improvement over the naive (random) approach and sometimes perform worse. Meanwhile, Augmentation refers to the approach that generates additional demonstrations and uses them along with original samples for ICL, for low-resource tasks (such as translation, reasoning, and classification) that do not contain enough samples to utilize the full capacity of LCLMs, showing substantial performance gains.
@@
This expansion of context length raises an important question: do previous sample selection strategies, designed for shorter context windows in earlier LMs, generalize to the many-shot ICL regime? To answer this, we systematically revisit existing sample selection strategies by conducting extensive experiments across 18 datasets spanning diverse tasks (namely, classification, translation, summarization, and reasoning) with multiple LCLMs. Our experiments include three types of sample selection methods: relevance, diversity, and difficulty-based sample selection, as outlined in Dong et al. (2023). From these experiments, we uncover novel and surprising findings: contrary to prevailing expectations that carefully selected ICL demonstrations would yield performance improvements, they are similarly effective with a simple random selection approach, offering no statistically meaningful improvements in almost all cases (Figure 1). An additional reason to prefer the naive sample selection approach is that it enables greater efficiency through key-value caching of in-context examples (as the same examples can be reused across multiple queries), unlike sophisticated sample selection methods where the examples vary for each sample.
@@
While the expanded context length in LCLMs allows us to focus less on selecting optimal subsets of examples, it introduces a new challenge: effectively utilizing this expanded capacity when the number of examples is limited. Specifically, in scenarios where available data is sparse (such as low-resource translation or reasoning tasks where annotated data samples are difficult or costly to obtain), the examples available only utilize a small fraction of the full context window. In other words, this mismatch between context capacity and example availability introduces a new direction in ICL research, shifting the focus from optimizing sample selection to maximally utilizing the long context window. To address this, we propose a simple yet effective data augmentation approach to increase the number of in-context examples, which consists of two consecutive steps: (1) generating synthetic examples and (2) filtering out low-quality examples via LCLM prompting prepended with randomly sampled real examples. Then, by adding these augmented data samples to the context, we significantly improve ICL performance.
@@
Moreover, we explore other key factors unique to LCLM-enabled ICL. Specifically, we investigate the capacity of LCLMs to comprehend extremely long context (where a large number of examples up to the context length are present), as well as how they handle scenarios in which some of these examples introduce noise. Through comprehensive analyses, we find that while performance generally improves as the number of in-context examples increases, it eventually plateaus and begins to decline as the context length approaches the limit. This diminishing return highlights the need to carefully balance context length and example quantity within the expanded capacity of LCLMs. In addition, we observe that LCLMs exhibit robustness to noisy examples in relatively simple tasks, but they become vulnerable to noise in more complex scenarios to which they might be less exposed during training, such as extremely low-resource translation tasks.
@@ 2.1 Background
A language model (LM), which takes an input sequence of tokens ${\bm{x}}=[x_{1},x_{2},\ldots,x_{n}]$ and generates an output sequence of tokens ${\bm{y}}=[y_{1},y_{2},\ldots,y_{m}]$, can be represented as follows: ${\bm{y}}=\texttt{LM}_{\theta}({\bm{x}})$, where $\theta$ is the set of model parameters that are typically fixed after training due to the high computational costs of fine-tuning. A long-context LM (LCLM) is an advanced LM Reid et al. (2024) that is designed to accommodate sequences with a large number of tokens (e.g., $n$ can exceed 1 million), typically far surpassing the context sizes of earlier LMs.
@@ 2.2 Experimental Setup
  • Reasoning: This task evaluates the capability of models to perform complex reasoning. We use three challenging datasets from Big Bench Hard (BBH) Suzgun et al. (2022) following the experimental setting of Long-Context Frontiers (LOFT) benchmark Lee et al. (2024a).
@@
  • Relevance: This method selects examples that are most similar to the input query to maximize the alignment of ICL demonstrations with the query. To compute semantic similarity between the query and each example, we use an embedding model Lee et al. (2024b).
  • Diversity: This method selects examples that are maximally distinct from each other to capture a broad coverage of features and characteristics within the task space. We first embed each example in a shared embedding space with Lee et al. (2024b) and utilize $k$-means clustering (where $k$ corresponds to the number of desired ICL examples) to group the examples into subcategories. We then select the example closest to each cluster center as the representative to capture a diverse subset of the task features.
  • Difficulty: This method selects examples based on their difficulty. We examine two approaches: the first method (called Curriculum) follows a curriculum learning paradigm where examples are ordered from easiest to hardest; the second one (called Hard) includes only difficult examples, as simpler examples may already be well-understood by models. To assess example difficulty, we use model-based evaluation Liu et al. (2023), which prompts LCLMs 30 times and averages difficulty scores weighted by probabilities.
LCLM Configurations for ICL
We consider LCLMs that support extensive token capacities to evaluate ICL performance in long-context, many-shot ICL scenarios. We focus on models that have context window lengths on the order of millions: Gemini 1.5 Flash, which can process up to 1 million tokens; Gemini 1.5 Pro, which can process up to 2 million tokens Reid et al. (2024). In addition, we also consider the Llama 3.1 70B model Dubey et al. (2024), which, while supporting the comparatively smaller context size of 128K tokens, is still considered an LCLM. For all experiments, we utilize the default hyperparameters for both Gemini and Llama. To provide a comprehensive view of performance under different shots, we vary the number of ICL examples, starting from one and sequentially doubling to 2, 4, 8, 16, 32, and so forth, until reaching either the context size limit or the maximum number of dataset samples, whichever is exhausted first. Furthermore, to ensure the reliability of our results, we conduct multiple runs for each experimental setup: 3 runs for translation and summarization tasks; 10 runs for reasoning and classification tasks. The prompts used to elicit responses from ICL are provided in Appendix A.
2.3 Experimental Results
@@
To see the performance of ICL with respect to the number of examples, we visualize results in Figure 3. Overall, for any sampling method, we observe that performance increases as the number of examples increases. Also, when the number of examples is relatively small, the relevance-based sample selection approach performs particularly well, as focusing on highly relevant examples maximizes learning effectiveness when using a small number on examples. However, as the number of examples increases, the performance gap between various sample selection methods diminishes, indicating that performance is less dependent on selection strategies in many-shot scenarios. Lastly, in the summarization task (where samples tend to be longer than those in other tasks), we observe an initial increase in performance as more examples are added, followed by a decline once the context becomes heavily populated with a large number of examples. We argue this decline likely reflects the challenges LCLMs face in processing extremely long contexts, and we offer more analysis and discussion in Section 4.2.
@@ 3.1 ICL Example Augmentation Approach
Formally, let $\mathcal{D}=\{({\bm{x}}_{i},{\bm{y}}_{i})\}_{i=1}^{k}$ be a dataset of available ICL examples for a given target task, where each example $({\bm{x}}_{i},{\bm{y}}_{i})$ represents an input-output pair. The objective is to generate a set of synthetic examples $\mathcal{D}^{\prime}=\{({\bm{x}}^{\prime}_{j},{\bm{y}}^{\prime}_{j})\}_{j=1}^{m}$ (to supplement the original dataset $\mathcal{D}$), such that the augmented set of examples $\mathcal{D}_{\texttt{AUG}}=\mathcal{D}\cup\mathcal{D}^{\prime}$ can increase the utilization of the available context capacity of LCLMs. To operationalize this, we generate each synthetic example (${\bm{x}}^{\prime}_{j},{\bm{y}}^{\prime}_{j}$) by prompting an LM with randomly selected real examples from $\mathcal{D}$ as context, to ensure that the generated data retains meaningful patterns and characteristics relevant to the task.
@@
Once the synthetic examples are generated, we filter out low-quality instances that may introduce noise or irrelevant information. To do this, we design a function $f$ that assigns a quality score to each synthetic example $({\bm{x}}^{\prime}_{j},{\bm{y}}^{\prime}_{j})$ based on its contextual relevance and alignment with real examples as well as overall quality. Specifically, each synthetic example is rated on a 5-point Likert scale by prompting the LM 30 times with the synthetic and 30 real examples. We then compute an aggregate score using a weighted average of scores with their corresponding probabilities from the LM. Only the synthetic examples that exceed a quality threshold, $\tau$, are retained in the augmented example set, as follows:
$\mathcal{D}_{\texttt{AUG}}=\mathcal{D}\cup\{({\bm{x}}^{\prime}_{j},{\bm{y}}^{\prime}_{j})\;|\;f({\bm{x}}^{\prime}_{j},{\bm{y}}^{\prime}_{j},\mathcal{D})\geq\tau\}_{j=1}^{m},$
where $f({\bm{x}}^{\prime}_{j},{\bm{y}}^{\prime}_{j},\mathcal{D})$ is the quality assessment function, and $\tau$ is the threshold value for filtering.
@@ 3.3 Experimental Results
As shown in Table 3, which compares the example augmentation approach (with random selection) to other sample selection strategies, the augmentation approach demonstrates substantial performance gains across various datasets, which can be attributed to the greater diversity and volume of ICL examples achieved through synthetic data generation, leading to the effective utilization of the context capacity of LCLMs. Also, like the random selection approach, our augmentation method allows the reuse of the same examples across all queries. Thus, due to key-value caching, the augmentation approach is as efficient as random selection while achieving superior performance.
Table 4: Results on ablation study, where w/o Filtering and w/o Original denote the ICL results based on augmented samples without filtering and without original samples, respectively. Only Original is the performance without generated samples.
@@
To see how each component in the augmentation approach contributes to performance gains, we conduct an ablation study. As shown in 3.3, we observe that the full augmentation method (called Augmentation), which uses both original examples and filtered synthetic examples in combination, achieves the best performance. In contrast, when the filtering step is omitted, performance decreases, indicating that filtering contributes positively by removing lower-quality synthetic examples. Also, a large performance drop occurs when original samples are excluded from the augmented set. This suggests that although filtering helps maintain quality, the synthetic samples generated still do not match the quality of the original examples. Thus, while our augmentation approach is effective, further research could improve data generation techniques to improve the quality of the synthetic examples.

### v3

@@ Abstract
In-Context Learning (ICL) is a technique by which language models make predictions based on examples provided in their input context. Previously, their context window size imposed a limit on the number of examples that can be shown, making example selection techniques crucial for identifying the maximally effective set of examples. However, the recent advent of Long Context Language Models (LCLMs) has significantly increased the number of examples that can be included in context, raising an important question of whether ICL performance in a many-shot regime is still sensitive to the method of sample selection. To answer this, we revisit these approaches in the context of LCLMs through extensive experiments on 18 datasets spanning 4 tasks. Surprisingly, we observe that sophisticated example selection techniques do not yield significant improvements over a simple random sample selection method. Instead, we discover that the advent of LCLMs has fundamentally shifted the challenge of ICL from that of selecting the most effective examples to that of collecting sufficient examples to fill the context window. Specifically, in certain datasets, including all available examples does not fully utilize the context window; however, by augmenting the examples in context with a simple data augmentation approach, we substantially improve ICL performance by 5%.
@@
Jinheon Baek1thanks: This work was conducted during an internship at Google.   Sun Jae Lee2   Prakhar Gupta2 Geunseob (GS) Oh2Siddharth Dalmia2Prateek Kolhar2 KAIST1    Google DeepMind2 jinheon.baek@kaist.ac.kr{sunjaelee, prakharguptaz, ohgs, pkolhar}@google.com
@@ 1 Introduction
Figure 1: Results of various sample selection approaches in 64-shot ICL with LCLMs. Approaches include Retrieval that selects examples similar to the target query, Diversity that aims for maximizing example variety, Curriculum that arranges examples in order from easiest to hardest, and Hard that uses only challenging examples, alongside Random that selects examples without any constraints. Results indicate that sample selection methods provide no significant improvement over the naive (random) approach and sometimes perform worse. Meanwhile, Augmentation refers to the approach that generates additional demonstrations and uses them along with original samples for ICL, particularly for low-resource tasks (such as translation, reasoning, and classification) that do not contain enough samples to utilize the full capacity of LCLMs, showing substantial performance gains.
@@
This expansion of context length raises an important question: do previous sample selection strategies, designed for shorter context windows in earlier LMs, generalize to the many-shot ICL regime? To answer this, we systematically revisit existing sample selection strategies by conducting extensive experiments across 18 datasets spanning diverse tasks (namely, classification, translation, summarization, and reasoning) with multiple LCLMs. Our experiments include multiple types of sample selection methods: relevance, diversity, and difficulty-based sample selection, as outlined in Dong et al. (2023). From these experiments, we uncover novel and surprising findings: contrary to prevailing expectations that carefully selected ICL demonstrations would yield performance improvements, they are similarly effective with a simple random selection approach, offering no statistically meaningful improvements in almost all cases (Figure 1). An additional reason to prefer the naive sample selection approach is that it enables greater efficiency through key-value caching of in-context examples (as the same examples can be reused across multiple queries), unlike sophisticated sample selection methods where the examples vary for each sample.
@@
While the expanded context length in LCLMs allows us to focus less on selecting optimal subsets of examples, it introduces a new challenge: effectively utilizing this expanded capacity when the number of examples is limited. Specifically, in scenarios where available data is sparse (such as low-resource translation or reasoning tasks where annotated data samples are difficult or costly to obtain), the examples available only utilize a small fraction of the full context window. This mismatch between context capacity and example availability introduces a new direction in ICL research, shifting the focus from optimizing sample selection to maximally utilizing the long context window. To address this, we propose a simple yet effective data augmentation approach to increase the number of in-context examples, which consists of two steps: (1) generating synthetic examples and (2) filtering out low-quality examples through LCLM prompting contextualized with real examples. Then, by adding these augmented data samples to the context, we significantly improve ICL performance.
@@
Moreover, we explore other key factors unique to LCLM-enabled ICL. Specifically, we investigate the capacity of LCLMs to comprehend extremely long context (where a large number of examples up to the context length are present), as well as how they handle scenarios in which some of these examples introduce noise. Through comprehensive analyses, we find that while performance generally improves as the number of in-context examples increases, it eventually plateaus and begins to decline as the context length approaches the limit. This diminishing return highlights the need to carefully balance context length and example quantity. Also, we observe that LCLMs exhibit robustness to noisy examples in relatively simple tasks, but become vulnerable to noise in more complex scenarios to which they might be less exposed during training, such as extremely low-resource translation tasks.
@@ 2.1 Background
Figure 2: Results of various sample selection approaches on ICL of 64 examples with LCLMs, where we average the performance over all models: Gemini Pro, Gemini Flash, and Llama 3.1, across four different tasks with 18 datasets. Each bar represents the averaged performance, with the upper and lower limits indicating standard deviation. See Figure 9 for results on each model.
@@
A language model (LM), which takes an input sequence of tokens ${\bm{x}}=[x_{1},x_{2},\ldots,x_{n}]$ and generates an output sequence of tokens ${\bm{y}}=[y_{1},y_{2},\ldots,y_{m}]$, can be denoted as follows: ${\bm{y}}=\texttt{LM}_{\theta}({\bm{x}})$, where $\theta$ is the set of model parameters. A long-context LM (LCLM) is an advanced LM Reid et al. (2024) that is designed to accommodate sequences with a large number of tokens (e.g., $n$ can exceed 1 million), typically far surpassing the context sizes of earlier LMs.
@@ 2.2 Experimental Setup
  • Reasoning: This task evaluates the ability of models on complex reasoning. We use four challenging datasets from Big Bench Hard Suzgun et al. (2022) following the experimental setting of Long-Context Frontiers (LOFT) benchmark Lee et al. (2024a), where each data sample follows a multiple-choice question answering format.
@@
  • Relevance: This method selects examples that are most similar to the input query to maximize the alignment of ICL demonstrations with the query. To compute semantic similarity between the query and each example, we use the state-of-the-art embedding model Lee et al. (2024b).
  • Diversity: This method selects examples that are maximally distinct from each other to capture a broad coverage of features within the task space. We embed each example in a shared embedding space with Lee et al. (2024b) and utilize $k$-means clustering (where $k$ corresponds to the number of desired ICL examples) to group the examples into subcategories. We then select the example closest to each cluster center as the representative to capture a diverse subset of the task features.
  • Difficulty: This method selects examples based on their difficulty. We examine two approaches: the first method (called Curriculum) follows a curriculum learning paradigm where examples are ordered from easiest to hardest; the second one (called Hard) includes only difficult examples, as simpler examples may already be well-understood by models. To assess example difficulty, we use model-based evaluation Liu et al. (2023) with the state-of-the-art LCLM Reid et al. (2024), which prompts a model 30 times and averages difficulty scores weighted by probabilities.
@@
LCLM Configurations for ICL
We consider LCLMs that support extensive token capacities to evaluate performance in long-context, many-shot ICL scenarios, such as those with context window lengths on the order of millions: Gemini 1.5 Flash (1M tokens) and Gemini 1.5 Pro (2M tokens) Reid et al. (2024). Also, we consider the Llama 3.1 70B model Dubey et al. (2024), which, while supporting the comparatively smaller context size of 128K tokens, is still considered an LCLM. To provide a comprehensive view of performance under different shots, we vary the number of ICL examples, starting from one and sequentially doubling to 2, 4, 8, 16, 32, and so forth, until reaching either the context size limit or the maximum number of dataset samples, whichever is exhausted first. Furthermore, to ensure the reliability of our results, we conduct multiple runs for each setup: 3 runs for translation and summarization tasks and 10 runs for reasoning and classification tasks. The prompts used to elicit responses from ICL are provided in Appendix A.
2.3 Experimental Results
@@
Table 2: Performance comparison of recent sample selection strategies (Auto-ICL, IDS, and ICCL) in many-shot ICL.
Methods Translation Summarization Reasoning Classification
Random 0.551 $\pm$ 0.005 0.311 $\pm$ 0.005 0.650 $\pm$ 0.020 0.539 $\pm$ 0.006
[…]
IDS 0.547 $\pm$ 0.003 0.313 $\pm$ 0.004 0.649 $\pm$ 0.018 0.537 $\pm$ 0.007
ICCL 0.553 $\pm$ 0.006 0.307 $\pm$ 0.006 0.653 $\pm$ 0.016 0.543 $\pm$ 0.006
Additional Results with Advanced Sample Selection Strategies
To further assess the robustness of our findings, we additionally evaluate several recent and more advanced sample selection strategies: Auto-ICL Yang et al. (2023), IDS Qin et al. (2023), and ICCL Liu et al. (2024), which have been proposed to improve ICL by selecting high-quality and relevant examples based on context or model feedback. As shown in Table 2.3, however, we find that none of these newer methods consistently outperform the simple random selection baseline across tasks, with performance fluctuations within the range of statistical variation. This reinforces our main claim that LCLMs are insensitive to the specific sample selection strategy in many-shot ICL.
Figure 4: Ratios of convex hull volume of in-context examples to the full dataset with varying numbers of ICL examples.
@@
To see the performance of ICL with respect to the number of examples, we visualize results in Figure 3. Overall, for any sampling method, we observe that performance increases as the number of examples increases. Also, when the number of examples is relatively small, the relevance-based sample selection approach performs particularly well, as focusing on highly relevant examples maximizes learning effectiveness with the limited number on examples. However, as the number of examples increases, the performance gap between various sample selection methods diminishes, indicating that performance is less dependent on selection strategies in many-shot scenarios. Lastly, in the summarization task (where samples tend to be longer than those in other tasks), we observe an initial increase in performance, followed by a decline once the context becomes heavily populated with a large number of examples. We argue this decline likely reflects the challenges LCLMs face in processing extremely long contexts, discussed in Section 4.2.
Analysis on Converge of ICL Examples
To further investigate why the performance gap between different approaches diminishes as the number of examples increases, we analyze the representational coverage of examples in-context relative to the full examples. Specifically, we measure the convex hull volume spanned by the embeddings of ICL examples (where we vary their numbers) and compare it to that of the entire dataset, which can serve as a proxy for how well the samples in-context capture the distribution of the full data. Our results, visualized in Figure 4, show that, when the number of ICL examples is moderate (e.g., 64), they already span over 80% of the convex hull volume of the full dataset in almost all tasks and datasets. This suggests that, beyond a certain threshold, adding more examples does not significantly improve coverage, as the selected examples, regardless of selection methods, can approximate the full data distribution.
@@ 3.1 ICL Example Augmentation Approach
Formally, let $\mathcal{D}=\{({\bm{x}}_{i},{\bm{y}}_{i})\}_{i=1}^{k}$ be a set of available ICL examples for a target task. The objective is to generate a set of synthetic examples $\mathcal{D}^{\prime}=\{({\bm{x}}^{\prime}_{j},{\bm{y}}^{\prime}_{j})\}_{j=1}^{m}$ (to supplement the original dataset $\mathcal{D}$), such that the augmented set of examples $\mathcal{D}_{\texttt{AUG}}=\mathcal{D}\cup\mathcal{D}^{\prime}$ can increase the utilization of the available context capacity of LCLMs. To operationalize this, we generate each synthetic example (${\bm{x}}^{\prime}_{j},{\bm{y}}^{\prime}_{j}$) by prompting an LM with randomly selected real examples from $\mathcal{D}$ as context, to ensure the generated data retains meaningful characteristics relevant to the task.
@@
Once the synthetic examples are generated, we filter out low-quality instances that may introduce noise or irrelevant information. To do this, we design a function $f$ that assigns a quality score to each synthetic example $({\bm{x}}^{\prime}_{j},{\bm{y}}^{\prime}_{j})$ based on its contextual relevance and alignment with real examples as well as overall quality. Specifically, each synthetic example is rated on a 5-point Likert scale by prompting the LM 30 times with the synthetic and 30 real examples. We then compute an aggregate score using a weighted average of scores with their corresponding probabilities from the LM. Only the synthetic examples that exceed the quality threshold, $\tau$, are retained in the augmented example set, as follows:
$\mathcal{D}_{\texttt{AUG}}=\mathcal{D}\cup\{({\bm{x}}^{\prime}_{j},{\bm{y}}^{\prime}_{j})\;|\;f({\bm{x}}^{\prime}_{j},{\bm{y}}^{\prime}_{j},\mathcal{D})\geq\tau\}_{j=1}^{m}.$
Notably, our data augmentation process is efficient, as it is performed offline and does not contribute to inference-time overhead. Also, it takes under 10 seconds per example, which can be done in parallel.
@@ 3.3 Experimental Results
As shown in Table 2.3, which compares the example augmentation approach (with random selection) to other sample selection strategies, the augmentation approach demonstrates substantial performance gains across various datasets, which can be attributed to the greater diversity and volume of ICL examples achieved through synthetic data generation, leading to the effective utilization of the context capacity of LCLMs. Also, like the random selection approach, our augmentation method allows the reuse of the same examples across all queries. Thus, due to key-value caching, the augmentation approach is as efficient as random selection while achieving superior performance.
Figure 5: Visualization of embedding-space with original and synthetic examples.
Table 5: Results on Similarity (embedding-level similarity between original and synthetic examples) and Volume (relative expansion of the convex hull with augmented examples).
Tasks Similarity Volume
Translation 0.5715 1.6563
Reasoning 0.8099 3.2328
Classification 0.6252 2.7931
Figure 6: Results with augmented examples according to the size of synthetic samples (Left) and context utilization of Gemini Pro (Right).
Table 6: Results on ablation study, where w/o Filtering and w/o Original denote results based on augmented samples without filtering and without original samples, respectively. Only Original shows results without augmentation.
@@
Analysis on Augmented Data
Beyond performance improvements, we analyze the characteristics of the augmented data to better understand its impact on ICL. First, as visualized in Figure 5, the embedding-space distribution of augmented examples closely follows that of real examples while expanding the overall data coverage, which suggests that the synthetic examples effectively capture task-relevant features without deviating substantially from the original data distribution. In addition, we further quantify this expansion through two metrics: the similarity between original and synthetic examples, and the relative expansion of the convex hull with augmented examples compared to that formed by original examples, and report results in Table 5. From this, we observe that while synthetic examples maintain a high degree of similarity to real examples (ensuring alignment with the task), they also significantly increase the volume of the data distribution. This balance between relevance and diversity highlights why our augmentation approach effectively enhances ICL performance.
Finally, we analyze the impact of the number of augmented examples on performance and their corresponding context utilization in LCLMs. As shown in Figure 3.3, while increasing the number of synthetic examples initially improves performance, it eventually plateaus, indicating diminishing returns. Also, despite augmentation improving context utilization, we find that even at peak performance, the augmented data occupies less than 3% of the full context capacity of LCLMs, which is significantly below the scale that LCLMs can handle (Figure 8). These suggest an interesting future work to develop more advanced augmentation strategies to increase the context utilization of LCLMs.
@@
To see how each component in the augmentation approach contributes to performance gains, we conduct an ablation study. As shown in Table 6, we observe that the full augmentation method (called Augmentation), which uses both original and filtered synthetic examples, achieves the best performance. In contrast, when the filtering step is omitted, performance decreases, indicating that filtering contributes positively by removing lower-quality examples. Also, a large performance drop occurs when original samples are excluded from the augmented set. This suggests that although filtering helps maintain quality, the synthetic samples generated still do not match the quality of the original examples. Thus, while our augmentation approach is effective, further research could improve data generation techniques to improve the quality of the synthetic examples.